wendy's mother is three times as old as wendy, and in 14 years, she will be twice as old as wendy will be then. How old will they be 7 years from now?

Answer by ankor@dixie-net.com(22740)

m = mother's present age
w = Wendy's present age
:
Write an equation for each statement
:
wendy's mother is three times as old as wendy,
m = 3w
and in 14 years, she will be twice as old as wendy will be then.
m + 14 = 2(w+14)
m + 14 = 2w + 28
m = 2w + 28 - 14
m = 2w + 14
replace m with (3w), from the 1st statement
3w = 2w + 14
3w - 2w = 14
w = 14 yrs is Wendy's present age
then
3(14) = 42 yrs is Mom's age
:
How old will they be 7 years from now? I'll let you figure this out